OpenStudy (anonymous):

When a coin is flipped, there are two possible outcomes: heads or tails. If a coin is flipped three times, what is the probability that there will be tails all three times? A. 1 over 8 B. 1 over 3 C. 1 over 2 D. 3 over 2

OpenStudy (anonymous):

whatever the outcome of three coin flips, the probability is always 1/2 * 1/2 * 1/2

OpenStudy (anonymous):

so its 1/2

OpenStudy (anonymous):

lol. do you think 2 * 2 * 2 = 2? =))))

OpenStudy (anonymous):

nope lol

OpenStudy (anonymous):

\[\huge \frac 12 \times \frac 12 \times \frac 12 = \frac{1\times 1\times 1}{2\times 2\times 2}\]